You were hired as a consultant to Delta State Laundry Services, whose target capital structure is 35% debt, 10% preferred, and 55% common equity. The interest rate on new debt is 6.5%, the yield on the preferred stock is 6%, the cost of retained earnings is 11.25%, and the tax rate is 40%. The firm will not be issuing any new stock. What is Delta State Laundry Services WACC?
Answer #1

The weighted average cost of capital is calculated using the below formula:

WACC=Wd*Kd(1-t)+Wps*Kps+We*Ke

where:

Wd= Percentage of debt in the capital structure.

Kd= The before tax cost of debt

Wps= Percentage of preferred stock in the capital structure

Kps=Cost of preferred stock

We=Percentage of equity in the capital structure

Ke= The cost of common equity.

T= Tax rate

WACC= 0.35*6.5%*(1 - 0.40) + 0.10*6% + 0.55*11.25%

     = 0.35*3.90% + 0.10*6% + 0.55*11.25%

= 1.3650% + 0.60% + 6.1875%

= 8.1525% 8.15%.

In case of any query, kindly comment on the solution.